OpenStudy (anonymous):

Five cards are dealt from a standard deck of 52 cards. a) How many different hands can be dealt? b) How many hands will contain exactly 4 hearts (and 1 other card)? c) How many hands will contain no face cards? d) How many hands will contain only spades or only diamonds? e) How many hands will contain only red cards? f) How many hands will contain no more than 2 hearts?

OpenStudy (anonymous):

Are you familiar with binomial coefficients? \[\left(\begin{matrix}N \\ k\end{matrix}\right) = \frac{N!}{(N-k)!k!}\]

OpenStudy (anonymous):

Uh no, sorry. I'm have grade 12 data management and I don't think I've learned that yet. I'm currently doing permutations and combinations

OpenStudy (anonymous):

It's a version of those :) The binomial coefficient tells us how many ways there are to pick k things from N things, without replacement, regardless of order. So: \[\left(\begin{matrix}52 \\ 5\end{matrix}\right) = \frac{52!}{(52-5)!5!} = \frac{52*51*50*...*3*2*1}{(47*46*...*3*2*1)*5*4*3*2*1} = \frac{52*51*50*49*48}{5*4*3*2*1}\] Tells us the number of ways to choose 5 cards from 52 (This is the total number of hands possible). This works rather simply. If we were choosing 52 cards from a deck of 52, and order mattered, we'd have 52! total combinations. We only want 5, so we remove 52-5=47 choices, leaving us at: \[\frac{52!}{(52-5)!} = \frac{52!}{47!}\] We then remove the stipulation or order mattering by dividing be the number of ways a given 5 cards could be arranged. That is, we divide by 5!, since we are drawing 5 cards. What we're left with is the binomial coefficient. If you're ok with using this, this problem is pretty easy. If you'd rather not, I'll have to think about how to do it a bit :)

OpenStudy (anonymous):

hmm. that seems confusing lol

OpenStudy (anonymous):

And just for clarification, the end of that long line got chopped off: \[\left(\begin{matrix}52 \\ 5\end{matrix}\right) = \frac{52*51*50*49*48}{5*4*3*2*1}\]

OpenStudy (anonymous):

Hmmm, without that, it's a long walk. Let's try, though. First, how many different hands of 5 cards can be made? For the first card, we can draw any one of the 52 cards. For the second card, we only have 51 to choose from. And it keeps decreasing for each card drawn. How many combinations could we make?

OpenStudy (anonymous):

311,875,200?

OpenStudy (anonymous):

i did 52x51x50x49x48

OpenStudy (anonymous):

Right, but that's if order matters. 52*51*50*49*48. We don't care if we get the Ace of Spades first, second, or last. We just care that we get it. Right? So we have to remove all the different combinations that give the same hand. Do you know how to do this?

OpenStudy (anonymous):

52c5?

OpenStudy (anonymous):

52 choose 5?!?! That's the binomial coefficient :D

OpenStudy (anonymous):

And yes, that's how you'd do it

OpenStudy (anonymous):

so, there are 2,598,960 different hands that can be dealt?

OpenStudy (anonymous):

Yup! Are you ok with using the 52c5 notation? If so, this just got a lot easier. \[52c5 = \left(\begin{matrix}52 \\ 5\end{matrix}\right)\] Which is what I was wanting to use :)

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

yeah, i guess i am okay with it now

OpenStudy (anonymous):

before we move on, can i ask you about something else?

OpenStudy (anonymous):

Excellent. So, you're done with a. Now, we want to insist on 4 cards being a heart, and 1 card that isn't a heart.

OpenStudy (anonymous):

Sure, whats up?

OpenStudy (anonymous):

(11+8 choose 2). do i add the 11 after or is it 19c2?

OpenStudy (anonymous):

this is unrelated to this question btw

OpenStudy (anonymous):

i just wanted clarification on this before moving on

OpenStudy (anonymous):

I'll do my best to answer it :)

OpenStudy (anonymous):

do u want the full question? i asked it last night.

OpenStudy (anonymous):

not sure if you can see it but here http://openstudy.com/users/dudehao#/updates/5362de59e4b02f72432c1ee6

OpenStudy (anonymous):

The chemists and technicians problem? Alrighty. What would you like clarified?

OpenStudy (anonymous):

the 11+8choose2. is it 19c2 or 8c2 then i add 11 to it

OpenStudy (anonymous):

With the notation that was being used, it would be 19c2 I'm checking to make sure I get the same answer for d, though.

OpenStudy (anonymous):

oh okay, thanks

OpenStudy (anonymous):

let me know once youre finished

OpenStudy (anonymous):

Yeah. And he did it in a much simpler way than what I started doing initially :)

OpenStudy (anonymous):

You do it by first insisting that you get at least 1 technician: 12c1 = 12 Then insist that you get at least 1 chemist: 9c1 = 9 The other two, we don't really care about. We just need 2 of them. There were 21 total people. We've chosen 2, so we have 19 left to choose from: 19c2 (12c1)(9c1)(19c2) = 12*9*(19c2)

OpenStudy (anonymous):

oh wow. thanks for clearing that up. now let's move on to the deck question

OpenStudy (anonymous):

Well, we'll approach this in the same way. First, we want to insist that we choose 4 hearts. There are 4 suits in a deck. So to specify that we want a heart, we do: (4c1) We then want 4 of these hearts. There are 13 hearts (13 cards of each suit) in a deck. Since we've already chosen our suit, we just pick 4 of them: (13c4) Then we just have to specify that the other card must come from another suit. How would we do this?

OpenStudy (anonymous):

uhm, 52c1?

OpenStudy (anonymous):

or 13c1?

OpenStudy (anonymous):

Close. We have to insist that it is NOT a heart, as well. 52c1 gives us a single random card from the deck. Could be a heart. Could not be. 13c1 gives us a single card from a suit, but it doesn't limit what that suit could be.

OpenStudy (anonymous):

39c1?

OpenStudy (anonymous):

That's one way to do it, sure :D Another way: Insist that the card is one of the other three suits: 3c1 Draw one from that suit: 13c1 Then we multiply it all together. (4c1)(13c4)(39c1) = ?

OpenStudy (anonymous):

I got 111,540

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Yes, thank you

OpenStudy (anonymous):

Now, how would you do the next one?

OpenStudy (anonymous):

how would i do c? what is a face card exactly?

OpenStudy (anonymous):

king queen and jack?

OpenStudy (anonymous):

Right. King, Queen, and Jack are face cards. You can use a choose method. There are also other methods. Try it out, let's see what you get :)

OpenStudy (anonymous):

for c, do I do 38c5?

OpenStudy (anonymous):

Why 38?

OpenStudy (anonymous):

because there are 14 facecards, i subtracted that from 52

OpenStudy (anonymous):

Jack Queen and King of each suit. 4 suits. 3*4 = 12 face cards

OpenStudy (anonymous):

so jokers dont count?

OpenStudy (anonymous):

Not in a 52 card deck. There are no jokers, unless they change the deck somehow. The cards are: Ace, 2, 3, 4, 5, 6, 7, 8, 9, 10, J, Q, K of each suit. 13 cards in each suit. 13*4 = 52

OpenStudy (anonymous):

ok i see, so it would be 36c5?

OpenStudy (anonymous):

52-12=40

OpenStudy (anonymous):

whoops true ahah

OpenStudy (anonymous):

So, 40c5 ways to make hands without face cards.

OpenStudy (anonymous):

yup, and i get 658,008

OpenStudy (anonymous):

Yep. What about the next part?

OpenStudy (anonymous):

26c5?

OpenStudy (anonymous):

only spades OR only diamonds. Not a combination of spades and diamonds.

OpenStudy (anonymous):

13c5?

OpenStudy (anonymous):

First, how would we decide how many hands could be made with only spades?

OpenStudy (anonymous):

misread the question whoops

OpenStudy (anonymous):

Ok, 13c5 will gives us a hand all from a single suit. So if we wanted this OR the same thing from another suit, we'd just add the probability to do it with the other suit to this. How many ways are there to make hands with only diamonds?

OpenStudy (anonymous):

it would also be 13c5?

OpenStudy (anonymous):

Yep. So (13c5)+(13c5) = 2*(13c5) = ?

OpenStudy (anonymous):

2574 is my answer

OpenStudy (anonymous):

Alright. And the next part?

OpenStudy (anonymous):

26c5

OpenStudy (anonymous):

Good :)

OpenStudy (anonymous):

thanks, you made this so much easier. yesterday i had no clue what I was doing lol

OpenStudy (anonymous):

Just a quick note: I made a mistake on part b The (4c1) shouldn't be there. We did a slightly different problem. So for b, you'd have: (13c4)(39c1)

OpenStudy (anonymous):

alright thanks and for the last part do i do 13c2?

OpenStudy (anonymous):

For the last part, we want NO MORE THAN 2 hearts. So, we can have no hearts, 1 heart, or 2 hearts. Find the number of ways to make a hand with no hearts. And the number of ways to make a hand with only 1 heart. And finally, with only 2 hearts. Then add them all up.

OpenStudy (anonymous):

39c5 + 13c1 + 13c2?

OpenStudy (anonymous):

Close. Remember, we have to draw 5 cards each time. 39c5 is fine, as you chose 5. 13c1, you only drew one card. We need to draw 4 more that are NOT hearts. 13c2, we need to draw 3 more.

OpenStudy (anonymous):

so 13c5 x2?

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

So, if I want only a single heart: I draw that heart (13c1) Then I draw 4 more cards that are NOT hearts. number of cards that aren't hearts: 52-13=39 So draw 4 that are not hearts: (39c4) So the number of ways to make a hand with only one heart is: (13c1)(39c4)

OpenStudy (anonymous):

so then the number of ways to make two hearts would be 13c2 x 39c3?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

so i get 795,093 in total

OpenStudy (anonymous):

I gotta head out now, but thank you for everything. I'll let you know if I need help with any more questions, thanks again!

OpenStudy (anonymous):

That's not what I get